OpenStudy (anonymous):

What are the solutions of 6x2 = 12x? x = 0, x = −2 x = 0, x = 2 x = 6, x = 2 x = 6, x = −2

OpenStudy (lgbasallote):

subtract 12 x from both sides 6x^2 - 12x = 0 factor out x x(6x - 12) = 0 does that help?

OpenStudy (anonymous):

a little

OpenStudy (lgbasallote):

wonderful

OpenStudy (anonymous):

For what values of x does it make the left hand side zero.

OpenStudy (anonymous):

All of the answer choices have x=0 as one. You can plug it to the equation @lgbasallote gave you and see how the left hand side becomes 0. Then you find another value for x for which this is true as well.

OpenStudy (anonymous):

6x(x-2)=0 x=0 or x=2

OpenStudy (anonymous):

x=0,x=2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome.......

OpenStudy (anonymous):

DUDE!!! why did you give the answer away?

OpenStudy (anonymous):

You didnt even show any work at all.

OpenStudy (anonymous):

talking to me...........?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Exactly if its so easy not to show your work then the asker should be able to answer it. If they still dont understand it then they can ask more questions

OpenStudy (anonymous):

whats to be detail in it....... i just took 6x common

OpenStudy (anonymous):

Thank you you understand.

OpenStudy (anonymous):

dude,,,,,,,,?my name is swati......